$SPAD/src/input d01alf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
e:EXPR FLOAT:=1/sqrt(abs(X-1/7))
f:ASP1(F):= retract e
a:SF:=0.0
b:SF:=1.0
points:Matrix SF:=[[1/7]]
npts:=ncols points
epsabs:SF:=0.0
epsrel:SF:=1.0e-4
liw:=max(npts+4,400)
lw:=max(2*npts+8,2*liw)
result:=d01alf(a,b,npts,points,epsabs,epsrel,lw,liw,-1,f)

———-

References^{[1] nothing}

4